OpenStudy (anonymous):

I REALLY DISLIKE INTEGRATION. Another one that has me stumped:

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1 }{ {1+4\sin^2x} }dx\]

OpenStudy (anonymous):

I've tried every simple substitution possible--it's gotta be some trig substitution, right?

OpenStudy (anonymous):

looks like arctangent or sommat

OpenStudy (anonymous):

how would you know D= how would I know

OpenStudy (anonymous):

because it is almost \[\frac{1}{1+x^2}\] only with x replaced by \(2\sin(x)\)

OpenStudy (anonymous):

i am trying to think of an easy way to get it wolfram tells me the answer, and gives me a method i do not like it does involve arctangent as i suspected

OpenStudy (anonymous):

oo, I tried making \[4\sin^2(x) = \tan^2(\theta)\] but it didn't work out for me. but that could just be because I'm working w/ the numbers wrong

OpenStudy (anonymous):

but the wolfram method is very complicated too complicated i think

OpenStudy (anonymous):

i can tell you what it does if you like

OpenStudy (anonymous):

I can see the worked solution using symbolab where they let x = arccot(u)

OpenStudy (anonymous):

=D the file name

OpenStudy (anonymous):

i considered just \(u=2\sin(x)\)

OpenStudy (anonymous):

Ahh, their solution looks like something I'd be expected to do Thank you for that

OpenStudy (anonymous):

integration like this is boring and stupid let the machine do it, it teaches you nothing

OpenStudy (anonymous):

really wish my prof thought so too